How long will it take for $2,800 to double if it is invested at 12% annual interest compounded 6 times a year? Enter in exact calculations or round to 3 decimal places.
a. How long will it take if the interest is compounded continuously?

Answer by Boreal(15235)

The first is 2800(1+(0.12/6))^6n=5600
or 2=1.02^6n
ln2=6n*ln(1.02)
6n=35.002
n=5.834 years
rule of 72 would say 6 years so reasonable result.
continuously is 2=e^rt, where P=Poe^rt, and P=2Po
so ln2=rt
ln2/0.12=t
t=5.776 years.